# Representing Molecules as Random Walks Over Interpretable Grammars

### v1

@@ Appendix F Generating Novel Random Walks
We illustration of the generation of the random walk with notation $G81\rightarrow G82\rightarrow G274\rightarrow G82:1$ in 15. Our graph resolves any ambiguity of whether to revisit $G82$ or attach a new copy of $G82$ to the molecule by attaching a colon for each newly attached motif that has a naming conflict. This is possible after augmenting $G$ with duplicates of the motif (see Appendix B), which, in practice, has a negligible increase the complexity of $G$.
Figure 15: Generation of a random walk $G81\rightarrow G82\rightarrow G274\rightarrow G82$; The possible transitions from $G81$, $G82$ and $G274$ are in $R$, $G$, $B$ (with thickness proportional to probability).
@@ 22                         M $\leftarrow$ attach(M, molecule(cur));
Our implementation in 7 handles the distinction between revisiting a previous node vs adding a new duplicate of the same motif as a previous node through mask__\__attach (new nodes which can be attached) vs mask__\__return (the node which the random walker can backtrack to). This distinction is done by creating duplicates of nodes for each revisit (see B).
For the case of molecules, we guarantee 100% validity since we can explicitly check the possible motifs which can be attached to $M$ at each step. When there are neither new motifs to attach, nor existing motifs to return to, the generation terminates with the current M being the final generation output. For the case of monomers, we cannot guarantee 100% validity because the random walk may not be closed (return to root__\__M). Please refer to the GitHub for details of the implementation.
Figure 16: Generation process of novel random walks on HOPV: $G262\rightarrow G305\rightarrow G181$ (Left) and $G239\rightarrow G297\rightarrow G202$ (Right)
As shown in figure 16, applying our generation method produces artifacts of learning that invite further scrutiny: “rules” of consecutive motifs. The second example in 16 shows there are only two possible motifs (green) that can be attached to the $G297$ end of a molecule with the $G239$ and $G297$ functional groups (ignoring the return back to G239, which transitions the state but does not attach a new motif). In the first example, the distribution of possible new motifs to attach to the $G305$ end of a molecule with $G262$ and $G305$ appears more uniform.
The termination criteria in 7 can be summarized as follows: see F.
Dataset Type Termination Criteria
Monomers root_M sampled again
Molecules Cannot sample node
Extracting Context-Sensitive Grammar Rules
One side product of our generation and verification procedure is the ability to extract “hard” rules. A hard rule is when a certain edge must be traversed (probability of 1) under a certain context and at a certain state. Although our context is invariant to the order of visited nodes thus far, we search for hard rules by using a best-first algorithm to store all promising trajectories.
Below is a compilation of hard rules learned by our model on the PTC dataset:
@@ Appendix G Detailed Case Study: Harvard Organic Photovoltaic Dataset
The Harvard Organic Photovoltaic Dataset (HOPV15) is a comprehensive collection that bridges experimental photovoltaic data with quantum-chemical calculations, serving as a crucial resource in the field of organic photovoltaics. This dataset includes experimental results from literature and corresponding quantum chemical data for a wide range of molecular conformers. These are analyzed using various density functionals and basis sets, including both generalized-gradient approximation and hybrid designs. A key feature of HOPV15 is its utility in calibrating quantum chemical results with experimental observations, aiding in the development of new semi-empirical methods, and benchmarking model chemistries for organic electronic applications. The dataset employs the Scharber model to compute the maximum percent conversion efficiencies for 350 studied molecules, focusing on their HOMO (Highest Occupied Molecular Orbital) values. Our segmentation approach involved systematically categorizing molecules based on their functional groups and ring structures. We separated standard functional groups (e.g., vinyl, alcohol) and individual rings (e.g., benzene, thiophene, pyrrole) to understand their unique contributions to photovoltaic properties. Additionally, we paid special attention to complex structures with consecutive rings, acknowledging their impact on the optical and electronic characteristics of the materials. These parameters are impactful to the molecular’s HOMO value, which are essential for calculating the open circuit potential and short circuit current density, leading to an understanding of percent conversion efficiency.
@@
Heuristic Based Fragmentation

### v2

@@ Appendix F Generating Novel Random Walks
We illustrate the generation of the random walk with notation $G81\rightarrow G82\rightarrow G274\rightarrow G82:1$ in Figure 14. Our graph resolves any ambiguity of whether to revisit $G82$ or attach a new copy of $G82$ to the molecule by attaching a colon for each newly attached motif that has a naming conflict. This is possible after augmenting $G$ with duplicates of the motif (see Appendix B.3), which, in practice, has a negligible increase the complexity of $G$.
Figure 14: Generation of a random walk $G81\rightarrow G82\rightarrow G274\rightarrow G82:1$; The possible transitions from $G81$, $G82$ and $G274$ are in Red, Green, and Blue (with thickness proportional to probability).
@@ 22                         M $\leftarrow$ attach(M, molecule(cur));
Our implementation in Algorithm 7 handles the distinction between revisiting a previous node vs adding a new duplicate of the same motif as a previous node through mask__\__attach (new nodes which can be attached) vs mask__\__return (the node which the random walker can backtrack to). This distinction is done by creating duplicates of nodes for each revisit (see B.3).
We guarantee 100% validity since we can explicitly check the possible motifs which can be attached to $M$ at each step. When there are neither new motifs to attach, nor existing motifs to return to, the generation terminates with the current M being the final generation output. Please refer to the GitHub for details of the implementation.
Figure 15: Generation process of novel random walks on HOPV: (Left) $G262\rightarrow G305\rightarrow G181$ and (Right) $G239\rightarrow G297\rightarrow G202$; The possible transitions from the first and second visited nodes are in Red and Green.
As shown in Figure 15, applying our generation method produces artifacts of learning that invite further scrutiny: “rules” of consecutive motifs. The second example in Figure 15 shows there are only two possible motifs (green) that can be attached to the $G297$ end of a molecule with the $G239$ and $G297$ functional groups (ignoring the return back to G239, which transitions the state but does not attach a new motif). In the first example, the distribution of possible new motifs to attach to the $G305$ end of a molecule with $G262$ and $G305$ appears more uniform.
F.1 Extracting Context-Sensitive Grammar Rules
One side product of our generation and verification procedure is the ability to extract “hard” rules. A hard rule is when a certain edge must be traversed (probability of 1) under a certain memory and at a certain state. Although our memory is invariant to the order of visited nodes thus far, we search for hard rules by using a best-first algorithm to store all promising trajectories. Table 9 is a compilation of hard rules learned by our model on the PTC dataset.
@@ Appendix G Detailed Case Study: Harvard Organic Photovoltaic Dataset
The Harvard Organic Photovoltaic Dataset (HOPV15) is a comprehensive collection that bridges experimental photovoltaic data with quantum-chemical calculations, serving as a crucial resource in the field of organic photovoltaics. This dataset includes experimental results from literature and corresponding quantum chemical data for a wide range of molecular conformers. These are analyzed using various density functionals and basis sets, including both generalized-gradient approximation and hybrid designs. A key feature of HOPV15 is its utility in calibrating quantum chemical results with experimental observations, aiding in the development of new semi-empirical methods, and benchmarking model chemistries for organic electronic applications. The dataset employs the Scharber model to compute the maximum percent conversion efficiencies for 350 studied molecules, focusing on their HOMO (Highest Occupied Molecular Orbital) values.
G.1 Segmentation Strategy
Our segmentation approach involved systematically categorizing molecules based on their functional groups and ring structures. We separated standard functional groups (e.g., vinyl, alcohol) and individual rings (e.g., benzene, thiophene, pyrrole) to understand their unique contributions to photovoltaic properties. Additionally, we paid special attention to complex structures with consecutive rings, acknowledging their impact on the optical and electronic characteristics of the materials. These parameters are impactful to the molecular’s HOMO value, which are essential for calculating the open circuit potential and short circuit current density, leading to an understanding of percent conversion efficiency.
@@ G.1 Segmentation Strategy
G.2 Heuristic Based Fragmentation
@@
G.3 Analysis of Learnt Representations
In this section, we perform an alternate and more accepted means of analysis than the 2D t-SNE analysis done in Section 4.4.2. We seek to understand the agreement between our property predictor’s learnt representation and the structural similarity over HOPV’s test set molecules. Since the final layer embedding is used for prediction, we expect molecules with similar properties to have similar embeddings.
In Figure 17, several groups of trends stand out. Particularly, highlighted in green are cases where the embedding similarity is high despite dissimilar HOMO property values; blue marks cases where the embedding similarity is low, and red marks sections that are similar in property, structure, and embedding. We detail each of these, basing comparison against molecule 50 for illustration:
  • For the topmost green section, molecules in the range 1-4 have similar components as those with higher HOMO values, though are much smaller in size and relatively disordered. For instance, molecules 3 and 4 each share key subcomponents (thiophene groups) with molecule 50, despite having quite different overall structure. The embedding similarity between (50, 4) and (50, 3) is thus medium-low and medium-high.
  • For the red sections along the diagonal, molecules in the ranges 14-16 and 17-26 cluster together. These tend to have an over-representation of electron-withdrawing groups in in non-symmetric locations in the structure, particularly methoxy, cyano, and carbonyl groups, without sufficient electron donating groups. Molecules 15 and 20 are shown as examples, and their embedding similarity is high. Blue outlines mark similar sub-groups between 15 and 20.
  • For the second-from-top green section, we again consider molecules in the range 18-26, where they show high similarity to the highest band in the range 47-63. These share many component structures, for instance thiophenes groups and derivatives. Molecule 23 is shown as an example, and has a barbituric acid core on one side, an electron withdrawing group, with methoxy groups on benzene rings on the other side, with a nitrogen atom between benzene rings, contributing to electron delocalization. The most likely explanation is that similar high-sterics groups have developed similar embeddings in this case.
  • For the bottom-right red section, molecules in the range 47-63 generally cluster together, reflecting the model’s ability to agree on both structural and property similarity. They tend to have an alternating pattern of electron-donating and electron-withdrawing groups which can increase the HOMO and provide a more direct pathway for charge transport. Yellow outlines mark matching and similar groups with molecule 50. In these cases, more than simply thiophene shows similar or the same structure. The embedding similarity between (50, 52), (50, 57), (52, 57) are all medium-high.
These insights show how complex molecule structure affects the measured property in this application, and how both structure and property are captured in the embedding. We hope the analysis provides more insights into how structural priors in our representation facilitates learning and generalization.
@@ G.3 Analysis of Learnt Representations
Figure 17: There are 64 molecules in this test set indexed from lowest to highest HOMO value. The above grid visualizes the distance between each pair of molecules as a cosine distance between the final layer embeddings of our model, with darker color representing lower distance (higher similarity). We use 4 quantiles, and refer to their ranges as low, medium-low, medium-high, and high similarity.